OpenStudy (anonymous):

what is the slope intercept of the line passing thourgh(-5,-2) and to the line y=1/2x-2

OpenStudy (anonymous):

Slope Intercept formula: y=mx+b

OpenStudy (anonymous):

ik that obviously but idk to do with the points and the equation

OpenStudy (anonymous):

Well if you obviously knew that then you would of known that the m is 1/2 and the b is -2. Then -5 is x and -2 is why. You have to put them in Slope intercept form.

OpenStudy (anonymous):

ok i got 8?

OpenStudy (anonymous):

i think your question is missing a word

OpenStudy (anonymous):

yea its parrel ot the line

OpenStudy (anonymous):

Alright. you see the slope of y=1/2x-2 is 1/2, right?

OpenStudy (anonymous):

wait? how is it 1/2?

OpenStudy (anonymous):

Slope intercept form y=mx+b m is the slope

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First use Point-slope form:\[y - y _{1}= m ( x - x _{1})\] If its parallel, it has the same slope, so m will be 1/2 The points will be (-5,-2), where Y1= -2 and X1= -5 Just plug these in, and isolate the y term on the left side to get slope-intercept form.

OpenStudy (anonymous):

o ok lol i did then but then i solved it. i wasn't supposed to solve it?

OpenStudy (anonymous):

nevermind did the math wrong

OpenStudy (anonymous):

Can you tell me what you get?

OpenStudy (anonymous):

i got -2=1/2+-2

OpenStudy (anonymous):

1/2x*

OpenStudy (anonymous):

You need to keep the x and y in there, and twhen you plug in in, it should look like: \[y + 2 = 1/2 ( x +5)\]

OpenStudy (anonymous):

isn't that point slope form?

OpenStudy (anonymous):

Yes, but in order to get to slope int form, you need to start here, and then change to slope int form

OpenStudy (anonymous):

Can you distribute the right side out?

OpenStudy (anonymous):

yes 1/2x+2.5

OpenStudy (anonymous):

Ok so now you have y + 2 = 1/2 x + 2.5 Now carry over the 2 to the right side

OpenStudy (anonymous):

y=1/2x+1/2

OpenStudy (anonymous):

Thats the equation in slope int form

OpenStudy (anonymous):

o ok thank you that really helped for my geom. final

OpenStudy (anonymous):

No problem